### v1

@@ 3.3 Multi-Concept Integration
3.3.1 Regional Customization Module.
Our key insight is to restrict the generation of each concept within a given region. As illustrated in the right part of Fig. 3, we propose the regional customization module in cross-attention layers, which makes it possible to integrate multiple LoRAs for multi-concept generation. Given a series of images $\mathcal{S}=\{X_{s}\}^{S}_{s=1}$ that represent $S$ concepts of interest, users can define the corresponding bounding boxes $B=\{b_{i}\}_{i=1}^{S}$ and region prompts $P_{r}=\{p_{i}\}_{i=1}^{S}$ for each concept, along with a base prompt $p_{base}$. The text embeddings $C=\{c_{i}\}_{i=1}^{S}$ of each concept can be acquired using the CLIP text encoder, calculated as follows:
$c_{i}=\tau_{\theta}(p_{i})+\tau_{\theta}(p_{base}),i=1,2,\cdots,S$ (7)
[…]
$K_{i}=W_{k}\cdot c_{i}+B_{i1}A_{i1}\cdot c_{i}$ (8)
$V_{i}=W_{v}\cdot c_{i}+B_{i2}A_{i2}\cdot c_{i}$ (9)
The attention operation is then applied to the query, key, and value vectors to derive the image feature as follows:
@@
$\operatorname{Attn}\left(Q_{i},K_{i},V_{i}\right)=\operatorname{Softmax}\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{d^{\prime}}}\right)V_{i}$ (10)
@@
where $W_{q}$, $W_{k}$, and $W_{v}$ serve as projection matrices for the query, key, and value inputs, respectively, and $d^{\prime}$ represents the output dimension of key and query features. The obtained image feature $\hat{f_{i}}=\operatorname{Attn}\left(Q_{i},K_{i},V_{i}\right)\in\mathbb{R}^{h_{i}\times w_{i}}$ retains its original dimensions. The final output is calculated as $\hat{F}[b_{i}]=\hat{f_{i}}$, where any overlapping regions have their related features $\hat{f_{i}}$ averaged.
The benefit of the proposed regional customization module is that each given region only interacts with the specific concept’s content during the cross-attention operation. This avoids the issue of mixing multi-concept features during cross-attention and allows each single-concept module to generate specific concepts following different text prompts in their respective regions. Moreover, the regional customization module brings minimal cost during inference, as evidenced in Tab. 2.
Figure 5: Qualitative comparisons. Our method outperforms all the compared methods significantly in terms of image fidelity and prompt alignment capability.
Table 2: Quantitative comparisons. Our approach surpasses other methods in all aspects of multi-concept customization. Both our training and inference time rank among the fastest. The inference time metric is determined by 50-step sampling. The top-performing method’s metrics are shown in red, while the second-best method’s metrics are in blue.
@@
Figure 6: CLIP-I v.s. Seg CLIP-I. Our Seg CLIP-I more accurately reflects the fidelity differences between the “good case” and the “bad case” compared to CLIP-I.
4 Experiment
4.1 Experimental Settings
4.1.1 Implementation Details.
All of our experiments are based on Stable Diffusion[23] v1.5 and are conducted on a single RTX3090. We set the rank of LoRA to be 16. We use the AdamW[11] optimizer with a learning rate of $8\times 10^{-5}$ and a batch size of 1, optimizing for 900 steps. During the inference stage, we utilize the DPM-Solver[12] for sampling 100 steps, with the guidance scale $\omega=7.5$.
4.1.2 Datasets.
Following Custom Diffusion[8], we conduct experiments on twelve subjects selected from the DreamBooth dataset[24] and CustomConcept101[8]. They cover a wide range of categories including two scene categories, two pets, and eight objects.
4.1.3 Evaluation Metrics.
As shown in Fig. 6, we have observed that when redundant elements dominate the calculation of the CLIP image score, it can result in high CLIP image scores for low-quality images. To address this issue, we suggest masking the objects in the source image that are truly relevant before computing the CLIP image score. We refer to this modified metric as Seg CLIP-I. As a result, we assess all the methods using three evaluation metrics: CLIP-I, Seg CLIP-I, and CLIP-T. (1) CLIP-I measures the average cosine similarity between the CLIP[21] embeddings of the generated images and the source images. (2) Seg CLIP-I is similar to CLIP-I, but all of the source images are processed with segmentation on the relevant objects. (3) CLIP-T calculates the average cosine similarity between the prompt CLIP embeddings and image CLIP embeddings.
4.1.4 Baselines.
We conduct comparisons between our method and four existing methods: Textual Inversion[1] (abbreviated as TI), DreamBooth[24] (abbreviated as DB), Custom Diffusion[8] (abbreviated as Custom), and Cones2[10]. To ensure consistency, we implement Textual Inversion, DreamBooth, and Custom Diffusion using their respective Diffusers[19] versions. For Cones2, we implement it using its official code. All experimental settings follow the official recommended settings of each method.
4.2 Qualitative Comparison
We validate our method and all the comparison methods on various prompts. Specifically, these prompts include illustrating the relative positions of concepts, generating concepts in new situations, creating multi-concept images with a specific style, and associating different concepts with different prompts. In Fig. 5, the results of all comparative methods are generated by the base prompt which is presented under the images, while the results of our method are jointly generated through both the base prompt and region prompts. These generated images demonstrate that our method excels in generating high-quality multi-concept images while effectively adhering to different prompts.
4.3 Quantitative Comparison
As presented in Tab. 2, our method demonstrates superior image alignment compared to other methods in the single-concept setting. Additionally, our method achieves comparable text alignment, showcasing its adaptability to various complex prompts. In the multi-concept setting, our method outperforms all the compared methods in the three selected metrics, notably excelling in CLIP-I and Seg CLIP-I. Moreover, with excellent image fidelity and prompt alignment ability, our method does not incur significant training and inference costs. This indicates the effectiveness of our efficient concept encoding and regional customization module regarding multi-concept customization.
@@ 4.4 Ablation Study
4.4.1 Without Regional Customization Module (Ours w/o region).
We conduct ablation experiments to validate the effectiveness of the regional customization module. In an attempt to directly load two LoRAs into the U-Net and generate images without utilizing the regional customization module, we find that
Table 3: Ablation Study.
Method CLIP-I Seg CLIP-I CLIP-T
Ours w/o Region 0.6918 0.7077 0.7104
Ours w/o QFormer 0.6911 0.6947 0.8238
Ours w/o ACN 0.6943 0.6951 0.8269
Ours 0.7135 0.7126 0.8380
this approach leads to feature confusion between different concepts and often only generates one object. As shown in Fig. 7, the features of the candle and teapot have fused to some extent under the prompt of “a $S^{*}$ candle and a $V^{*}$ teapot with a city in the background”, resulting in an output that contains both candle and teapot features. Moreover, such a generation method without the regional customization module produces only one object most of the time, which significantly lowers its CLIP-T in comparison to our approach as presented in Tab. 3.
Figure 7: Visualization of Ablation Studies. We visualize the impact of removing specific components from our framework individually to demonstrate the importance of these methods.
4.4.2 Without Training QFormer (Ours w/o QFormer).
In MiniGPT4[35], the QFormer is completely frozen, and a linear layer is trained to project the output of the QFormer. Following this, we experiment by freezing the QFormer and training only a linear layer. As illustrated in Tab. 3, this approach results in a decrease in the learning capacity of concepts and consequently a decline in the representation capability of the generated customized embeddings. As a result, there is a noticeable decrease in all three selected metrics. Images generated by this approach also have lower fidelity compared to our method as shown in Fig. 7.
4.4.3 Without adaptive concept normalization (Ours w/o ACN).
We conduct an ablation experiment to verify the effectiveness of the Adaptive Concept Normalization (ACN), which is used to mitigate the domain gap between the customized embedding and the other word embedding. As illustrated in Fig. 7 and Tab. 3, this domain gap leads to a decline in the customization ability, resulting in a notable decrease in image fidelity and prompt alignment ability compared to our method.
Table 4: User Study. The numbers represent the percentage (%) of users favoring the results of the methods. Our method is favored by the majority of users in the multi-concept setting for both text and image alignment.
TI[1] DB[24] Custom[8] Cones2[10] Ours
Text Alignment Image Alignment Text Alignment Image Alignment Text Alignment Image Alignment Text Alignment Image Alignment Text Alignment Image Alignment
Single-Concept 37.72 39.20 45.22 58.08 65.63 22.48 45.58 47.94 63.01 69.17
Multi-Concept 21.13 34.02 38.00 23.96 35.50 18.64 44.93 38.04 65.64 62.64
4.5 User Study
We conduct user studies for all methods. For text alignment, we display an image of each method alongside its prompt and inquire, “Which images accurately represent the text description?”. For image alignment, we exhibit various training and generated images and ask, “Which images align best with the target images?”. Each questionnaire comprises 50 such inquiries. From a total of 460 received questionnaires, 72 are deemed invalid, resulting in 388 valid responses. As shown in Tab. 4, users prefer our approach for multi-concept generation regarding both text and image alignment.
Figure 8: Challenge cases. Our method consistently generates high-fidelity images when customizing a larger number of concepts.
4.6 Challenge Cases
@@
Our method theoretically enables the combination of an unlimited number of concepts, facilitating true multi-concept customization. In Fig. 8, we present the results of customizing three and four concepts, including a total of 10 concepts. The visualization demonstrates that our method consistently generates high-quality customization images when facing a larger number of concepts, further demonstrating the superiority of our approach.
Figure 9: Extensions of our proposed MultiBooth. Our methods can also be applied to LoRA-based DreamBooth and ControlNet.
4.7 Extensions
4.7.1 Enable DreamBooth to Perform Multi-concept Generation.
Our regional customization module can guide multiple LoRAs to conduct multi-concept generation, and it is adaptable to any LoRA-related method. An illustration of transferring the regional customization module to the LoRA-based DreamBooth[24] method has been provided. As shown in Fig. 9, with the support of the regional customization module, DreamBooth[24] is capable of the generation of multiple concepts. However, the absence of customized embeddings results in a decrease in fidelity.
4.7.2 Apply ControlNet to our MultiBooth.
Our method is compatible with ControlNet[34] to achieve structure-controlled multi-concept generation. As illustrated in Fig. 9, our model inherits the architecture of the original U-Net model, resulting in satisfactory generations through seamless integration with pre-trained ControlNet without additional training.
@@ 5 Conclusion
In this paper, we introduce MultiBooth, a novel and efficient framework for multi-concept customization(MCC). Compared with existing MCC methods, our MultiBooth allows plug-and-play multi-concept generation with high image fidelity while bringing minimal cost during training and inference. By conducting qualitative and quantitative experiments, we robustly demonstrate our superiority over state-of-the-art methods within diverse multi-subject customization scenarios. Since current methods still require training to learn new concepts, in the future, we will investigate the task of training-free multi-concept customization based on our MultiBooth.

### v2

@@ 3.3 Multi-Concept Integration
Regional Customization Module.
To integrate multiple single-concept modules for multi-concept generation, we propose the Regional Customization Module (RCM) in cross-attention layers. The key insight of our RCM is to generate each concept within the specified region and allow different concepts to interact accurately in overlapping regions.
As shown in the right part of Fig. 2, given a base prompt $p_{base}$ describing the desired generated results, we can obtain the bounding boxes $B=\{b_{i}\}_{i=1}^{S}$ and the corresponding region prompts $P_{r}=\{p_{i}\}_{i=1}^{S}$ for each concept either through user-defined methods or automated processes (see Section 4.3). The region prompt guides the concept generation within each specific region, while the base prompt ensures interaction among concepts across different regions. As a result, the text embeddings $C=\{c_{i}\}_{i=1}^{S}$ for each region can be acquired through the combination of the region prompt and the base prompt:
$c_{i}=\tau_{\theta}(p_{i})+\tau_{\theta}(p_{base}),i=1,2,\cdots,S,$ (5)
where $c_{i}\in\mathbb{R}^{k\times d}$, $k$ is the the maximum length of input words and $\tau_{\theta}(\cdot)$ is the CLIP text encoder.
Then, we integrate the text guidance from text embeddings and the concept information in LoRA into each region simultaneously within the cross-attention layers. As shown in Fig. 3, the image feature $F\in\mathbb{R}^{h\times w}$ is the input of RCM. For the $i^{th}$ concept, the image feature $F$ is cropped using the bounding box $b_{i}\in\mathbb{R}^{h_{i}\times w_{i}}$, resulting in the partial image feature $f_{i}\in\mathbb{R}^{h_{i}\times w_{i}}$. With $f_{i}$, we can obtain the query vector $Q_{i}$ through $Q_{i}=W_{q}\cdot f_{i}$. Next, we derive the key and value vector $K_{i}$ and $V_{i}$ using the text embedding $c_{i}$ and corresponding LoRA parameters $\{A_{ij},B_{ij}\}_{i=1}^{S}$ through:
$K_{i}=W_{k}\cdot c_{i}+B_{i1}A_{i1}\cdot c_{i},$ (6)
$V_{i}=W_{v}\cdot c_{i}+B_{i2}A_{i2}\cdot c_{i},$ (7)
where $A_{ij}\in\mathbb{R}^{r\times k}$ and $B_{ij}\in\mathbb{R}^{d\times r}$, $j=1$ and $j=2$ indicating the low-rank decomposition of $W_{k}$ and $W_{v}$ respectively. In order to derive the text-aligned image feature with concept information, we then apply the attention operation to the query, key, and value vectors:
@@
$\operatorname{Attn}\left(Q_{i},K_{i},V_{i}\right)=\operatorname{Softmax}\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{d^{\prime}}}\right)V_{i},$ (8)
@@
where $d^{\prime}$ represents the output dimension of key and query features. The image feature $\hat{f_{i}}=\operatorname{Attn}\left(Q_{i},K_{i},V_{i}\right)\in\mathbb{R}^{h_{i}\times w_{i}}$ contains both the text guidance and concept information through the attention mechanism and retains its original dimensions. For overlapping regions, we use a weighted average strategy to ensure the generation of each concept:
$\hat{f}=\frac{1}{\eta}\cdot\sum_{i=1}^{\eta}{w_{i}\cdot\hat{f}_{i}},\ \sum_{i=1}^{\eta}{w_{i}}=1,\ \bigcup_{i=1}^{\eta}{b_{i}}\neq\varnothing,$ (9)
where $\eta$ is the number of overlapping concepts, $\hat{f}$ is the output feature of the overlapping region, $w_{i}$ is the average weight of the $i^{th}$ concept. The setting of $w_{i}$ is further discussed in $Suppl$.
Compared to (Kumari et al. 2023; Liu et al. 2023), our RCM offers more flexible and precise customization that cannot be achieved solely through text prompts. Once the single-concept modules are obtained, RCM can combine multiple single-concept modules in a plug-and-play manner to perform multi-concept generation without retraining. With bounding boxes indicating the regions of the generated concepts, RCM can generate each concept according to different region prompts (see Section 4.1) and handle complex object interactions under the guidance of the base prompt (see Section 4.3). Moreover, despite the superior multi-concept customization performance achieved by our RCM, it incurs minimal cost during inference. This is because the RCM generates all the customized concepts simultaneously, rather than sequentially, which is further discussed in Section 4.3. We also provide a thorough comparison between our RCM and other layout T2I methods (Lian et al. 2023; Xie et al. 2023), detailed in Section 4.2.
4 Experiment
Implementation details. All of our experiments are based on Stable Diffusion v1.5 and are conducted on one RTX3090. We set the rank of LoRA to be 16. During training, we randomly select text prompts $P_{s}$ from the CLIP ImageNet templates (Radford et al. 2021) following the Textual Inversion (Gal et al. 2022). During training, we optimize for 900 steps with a learning rate of $8\times 10^{-5}$. During inference, we sample for 100 steps with the guidance scale $\omega=7.5$. More detailed settings can be found in the $Suppl$.
Datasets. Following Custom Diffusion (Kumari et al. 2023), we conduct experiments on twelve subjects selected from the DreamBooth dataset (Ruiz et al. 2023) and CustomConcept101 (Kumari et al. 2023). They cover a wide range of categories including two scene categories, two pets, and eight objects.
4.1 Comparative Study
We conduct comparisons between our method and four existing methods: Textual Inversion (TI) (Gal et al. 2022), DreamBooth (DB) (Ruiz et al. 2023), Custom Diffusion (CD) (Kumari et al. 2023), and Cones2 (Liu et al. 2023).
Qualitative comparison. As shown in Fig. 4, TI and DB are limited to generating a single concept, whereas CD and Cones2 can produce multiple concepts but struggle with maintaining high fidelity. In contrast, our method excels in multi-concept generation, achieving both high image fidelity and prompt alignment, even in challenging long-format scenarios (third and fourth rows).
Figure 5: Qualitative ablation results.
Quantitative comparison. We assess all the methods using three evaluation metrics: CLIP-I, Seg CLIP-I, and CLIP-T. (1) CLIP-I measures the average cosine similarity between the CLIP (Radford et al. 2021) embeddings of the generated images and the source images. (2) Seg CLIP-I is similar to CLIP-I, but all the subjects in source images are segmented. (3) CLIP-T calculates the average cosine similarity between the embeddings of prompt and image. As presented in Tab. 3, our method demonstrates superior image alignment and comparable text alignment in the single-concept setting. In the multi-concept setting, our method outperforms all the compared methods in the three selected metrics. Moreover, with excellent image fidelity and prompt alignment ability, our method does not incur significant training and inference costs.
Table 2: Quantitative comparisons. The best and second best results are in red and blue, respectively.
@@ 4.1 Comparative Study
Table 3: Quantitative ablation results.
Method CLIP-I Seg CLIP-I CLIP-T
w/o Region 0.691 0.707 0.710
w/o QFormer 0.691 0.694 0.823
w/o ACN 0.694 0.695 0.826
Ours 0.713 0.712 0.838
@@ 4.2 Ablation Study
Regional Customization Module (RCM). We first verify the effectiveness of RCM by simply removing it. As shown in Fig. 5 and Tab. 3, without RCM, the features of the candle and teapot have fused to some extent. To further validate the effectiveness of RCM, we retain our single concept learning (SCL) and replace our RCM with other layout T2I methods. As shown in the right of Fig. 6, we select two representative methods: LLM-grounded Diffusion (LG) (Lian et al. 2023) and BoxDiff (Xie et al. 2023), with the bounding boxes used displayed on the left. On the one hand, LG (Lian et al. 2023) denoises each concept within the bounding boxes sequentially and then integrates them at the latent level, resulting in concept fusion in the overlapping regions. On the other hand, BoxDiff (Xie et al. 2023) employs the cross-attention map to construct a loss function for updating the latent variables. Although it can generate two concepts simultaneously, it suffers from low image fidelity. Furthermore, neither of these methods can handle complex object interactions according to the given text prompt. In contrast, our method allows different single-concept modules to target specific regions at the cross-attention level, thereby generating multiple concepts simultaneously. By using a base prompt to guide complex object interactions across various regions, we can produce images with both high image fidelity and precise text alignment.
QFormer and Adaptive Concept Normalization (ACN). We also demonstrate the effectiveness of the QFormer and the ACN by removing them either. As shown in Fig. 5 and Tab. 3, without QFormer or ACN, although our method can generate multiple concepts, their fidelity has decreased. In contrast, our full method can faithfully perform multi-concept generation.
4.3 Discussions
How to handle complex object interaction? Based on Eq. 5, each region is guided by both a regional prompt and a global base prompt. For overlapping regions, the base prompt directly guides their generation along with the regional prompts, ensuring accurate interaction and mitigating concept mixture. As shown in the left of Fig. 6, the absence of the base prompt leads to failures in object interactions or concept feature confusion. However, with the guidance of the base prompt, our method can effectively manage complex object interactions within the overlapping region.
Figure 6: Ablation study on the base prompt and RCM.
Inference time $\times$$N$ for $N$ concepts? We also analyze the inference time of our method with the increasing number of concepts. As shown in Tab. 4, the inference time of our method increases only slightly as the number of concepts grows. This is because increasing concepts only leads to additional cross-attention computation in our RCM; other operations, like self-attention, residual addition, etc. remain the same as generating a single concept.
Table 4: Inference Time with more concepts.
2 Concepts 3 Concepts 4 Concepts
Inference Time 8.29s 10.07s 10.53s
4.4 Further Analysis
@@
Customizing more than 2 concepts. Our method theoretically enables the combination of an unlimited number of concepts, facilitating true multi-concept customization. In Fig. 7, we present the results of customizing three and four concepts. The visualization demonstrates that our method consistently generates high-quality customization images when facing a larger number of concepts, further validating the superiority of our approach.
Figure 7: Challenge cases.
Extensions. Our approach is compatible with ControlNet (Zhang, Rao, and Agrawala 2023) for structure-controlled multi-concept generation. Besides, the RCM can guide multiple LoRAs in conducting multi-concept generation and is adaptable to any LoRA-related method, including the LoRA-based DreamBooth (Ruiz et al. 2023). Visualization results are presented in Fig. 8.
Figure 8: Visualization results of extensions.
@@ 5 Conclusion
We introduce MultiBooth, a novel and efficient framework for multi-concept customization (MCC). Compared with existing MCC methods, our MultiBooth allows plug-and-play multi-concept generation with high image fidelity while bringing minimal cost during training and inference. By conducting qualitative and quantitative experiments, we demonstrate our superiority over state-of-the-art methods within diverse customization scenarios. We believe that our approach provides a novel insight for the community.
6 Acknowledgments
This work was supported by the STI 2030-Major Projects under Grant 2021ZD0201404.